# Kishu: Time-Traveling for Computational Notebooks

### v3

@@ 5.1. Incremental Checkpointing
The Checkpoint Graph is a directed tree of (incremental) checkpoints representing the branch-based state evolution. Nodes are added for each checkpoint Kishu performs, and timestamped according to the completion time $t$ of the corresponding cell execution (we refer to the timestamped node and cell execution as node $t$ and CE $t$, respectively, for simplicity). The Checkpoint Graph maintains a head node, which tracks the user’s current state. Each node $t$ contains the state delta consisting of (only) Co-variables updated by CE $t$. Co-variables stored in the Checkpoint Graph are versioned according to their corresponding CE $t$:
@@
Versioned Co-variables are analogous to versioned datasets: the same Co-variable (w.r.t. variable membership, i.e., $\mathcal{X}=\{x_{1},...,x_{i}\}$) can take on multiple values throughout a notebook session being updated by different cell executions. LABEL:fig:checkpoint_graph show an example: CE $t_{3}$ creates the Co-variable {plot}, hence it is stored in node $t_{3}$ (red) as the Versioned Co-variable $((plot),t_{3})$.
@@
Kishu creates a new node $t$ in the Checkpoint Graph after each CE $t$ containing (1) the versioned Co-variables in the state delta of CE $t$, (2) the code of CE $t$, and (3) the versioned Co-variables (possibly stored in previous checkpoints) accessed by CE $t$ (§ 4.3). For example, node $t_{3}$ in LABEL:fig:checkpoint_graph (blue) contains the code of CE $t_{3}$ (“plot=gmm.result()“) and its dependency on the versioned Co-variable from node $t_{2}$ $((gmm),t_{2})$ (dashed line). Notably, the state delta, code, and variable accesses are respectively analogous to the update, operation, and dependencies in database logging and versioning. The new node $t$ is written under the head node $s$, and and a parent-child relationship is added from the head node $s$ to the new node $t$ (now the new head node).
@@
If Kishu cannot write an updated Co-variable in the state delta into the Checkpoint Graph (e.g., it contains an unserializable object such as a generator (Foundation, 2023b) or hash (Foundation, 2023c)), Kishu simply skips its storage. Instead, upon checkout, the missing (unserializable) Co-variable will be restored through fallback recomputation enabled by the cell code and dependencies stored in the Checkpoint Graph node, which we discuss in § 5.3.
@@ 5.2. Efficient State Restoration
When (incremental) checkout is requested, Kishu aims to accurately restore the current state to the target state in the fastest manner possible. To do so, it must identify the contents of the target state according to its timestamp, analogous to how timestamped state snapshots are defined in MVCC (Bernstein and Goodman, 1983); instead of versioned tables, we identify Versioned Co-variables in the (timestamped) target state:
@@
The Session State [^5] Kishu stores snapshots of Session State metadata (i.e., references to contained Co-variables) in Checkpoint Graph nodes. at timestamp $t$ is a set of $n$ Versioned Co-variables $\{(\mathcal{X}_{1},t_{1}),...,(\mathcal{X}_{n},t_{n})\}$ such that for each $(\mathcal{X}_{j},t_{j})$, $1\leq i\leq n$:
@@
    $t_{j}$ is an ancestor of $t$ on the Checkpoint Graph.
@@
    There must not exist another versioned Co-variable $(\mathcal{Y}_{k},t_{k})$ such that $\mathcal{X}_{j}\cup\mathcal{Y}_{k}\neq\emptyset$ and $t_{k}$ is a child of $t_{j}$ and ancestor of $t$.
The session state at timestamp $t$ (state $t$ for brevity) is the set of all Versioned Co-variables that are (1) available in the namespace after CE $t$ and (2) have not been overwritten by a newer Versioned Co-variable prior to CE $t$. For example, in LABEL:fig:restore_diff, state $t_{3}$ (top-left) consists of Versioned Co-variables ($(\{plot\},t_{3})$, ($(\{gmm\},t_{2})$, and ($(\{df\},t_{1})$. It does not contain ($(\{gmm\},t_{1})$ as it was overwritten by CE $t_{2}$ (gmm.fit(k=3)) which writes ($(\{gmm\},t_{2})$. Each state $t$ dictates which Versioned Co-variables should be loaded from various nodes on the Checkpoint Graph for checkouts; to efficiently accomplish (incremental) checkout, Kishu identifies the difference between the current and target states in terms of the (versioned) Co-variables that need to be updated: some Co-variables do not need to be updated when converting the current state to the target state. They can be identified via the Checkpoint Graph:
@@
A Co-variable $\mathcal{X}$ is identical between current and target states $t_{a}$ and $t_{b}$ if its versioned counterpart is consistent across $t_{a}$, $t_{b}$, and $t_{c}$, i.e., no cell execution between (1) node $t_{a}$ and node $t_{c}$ and (2) node $t_{b}$ and node $t_{c}$ updated $\mathcal{X}$, hence does not need updating when checking out from $t_{a}$ to $t_{b}$. For example, in LABEL:fig:restore_diff, if checking out from current state $t_{5}$ to target state $t_{3}$, the Co-variable {df} (blue) is identical between the states as no cell execution between (1) node $t_{1}$ and node $t_{3}$ and (2) node $t_{1}$ and node $t_{5}$ updated it.
Otherwise, if the Co-variable $\mathcal{X}$ has diverged between the current and target session states, it will need to be updated (by either loading the appropriate Versioned Co-variable or deleting it) to complete the checkout to the target state. For example, the Co-variable {gmm} (red) has diverged between nodes $t_{5}$ and $t_{3}$ as their parents ($t_{4}$ and $t_{2}$) both updated gmm with their cell execution (fitting with k=3 and k=10), hence, gmm (and plot) needs to be updated via loading ({gmm},$t_{2}$) if checking out from state $t_{5}$ to state $t_{3}$.
@@ 5.3. Robust Restoration
As each Checkpoint Graph node $t$ contains the code of CE $t$ and (2) which Versioned Co-variables $(\mathcal{X}_{j},t_{j})$ (stored in previous deltas, $t_{j}<t$) CE $t$ accessed (§ 5.1), any Versioned Co-variable $(\mathcal{X},t)$ in the state delta of node $t$ can be recomputed by (1) loading the accessed Versioned Co-variables from previous checkpoints into a temporary namespace, then re-running CE $t$. For example, in LABEL:fig:restore_fallback, suppose Versioned Co-variable ({plot}, $t_{3}$) (green) fails to load when checking out to session state $t_{3}$. The versioned Co-variable (({gmm}, $t_{2}$) is required to rerun CE $t_{3}$ (red); therefore, ({gmm}, $t_{2}$) is loaded from the parent node $t_{2}$, and after rerunning CE $t_{3}$ on the input ({gmm}, $t_{2}$), ({plot}, $t_{3}$) is restored.
@@
Kishu’s fallback recomputation is dynamic and recursive—if another Co-variable is missing or fails to load when retrieving the inputs (i.e., accessed Co-variables) for performing fallback recomputation, fallback recomputation can be recursively performed for that Co-variable. For example, if ({gmm}, $t_{2}$) from node $t_{2}$ fails to load (as part of fallback recomputation for ({plot}, $t_{3}$) from node $t_{3}$), it itself can be recomputed by loading ({gmm}, $t_{1}$) from node $t_{1}$ and rerunning CE $t_{2}$ (blue).
@@
Kishu guarantees exact restoration for all serializable Co-variables during checkout, i.e., they will have the same bytestring representation before and after checkout if there are no hidden serialization errors (§ 6.2). While Kishu is capable of restoring problematic Co-variables via fallback recomputation, it currently does not support exactly restoring Co-variables that both (1) fail to store or load and (2) are created through non-deterministic means (e.g., random generators). This is a limitation similarly present in Spark (Zaharia et al., 2010) and Ray’s (Moritz et al., 2018) lineage-based fault tolerance; however, in our case, unserializable objects are sufficiently rare in data science libraries (§ 7.2), hence we consider this limitation to be acceptable.
@@ 6.1. Implementation
Currently, Kishu uses SQLite (SQLite, 2024) to store Versioned Co-variables in the Checkpoint Graph. However, any storage mechanism can be used in its place—even in-memory ones if the user wants to maximize checkpointing/checkout efficiency.
@@ 6.2. Design Considerations
Certain object classes may contain incorrect serialization instructions, which, despite being able to be stored/loaded to/from storage, result in silent errors. Kishu currently assumes that the serialization instructions are correctly implemented for all objects in terms of equality before and after pickling (i.e., x == unpickle(pickle(x)) and does not prevent these silent errors. However, these cases are rare (§ 7.2.1), and Kishu provides a blocklist file for users to force fallback recomputation for Co-variables containing objects belonging to these classes.
@@
While Kishu’s VarGraphs can generalizably detect Co-Variable updates, there are specific cases in which more efficient detection methods are applicable, for example, (1) hashing (e.g., XXH64 (xxHash, 2023)) for array-like data (Team, 2024a)) and (2) rule-based static cell (e.g., df.head) identification for skipping update detection (§ 2.2). Kishu currently uses hash-based detection for some common array-likes (e.g., TensorFlow’s tensors (TensorFlow, 2024a)), but can be extended to incorporate (1) other efficient detection methods for specific data classes and (2) rule-based cell analyses.
@@ 7. Experimental Evaluation
    Low Checkpoint Times: Kishu’s incremental per-cell-execution checkpoints are created in up to 5.12$\times$ less time compared to the next best mechanism (§ 7.4)
@@ 7.1. Experiment Setup
We select 8 data science notebooks from Kaggle Grandmaster-level users or Github-hosted tool tutorials (e.g., Ray) (LABEL:tbl:workload), each featuring a popular data science library, which we categorize as in-progress (3/8) or final (5/8), with the former containing out-of-order cell executions and the latter lacking them. Notably, the final notebooks are also cleaned and contain memory/runtime optimizations. We empirically verify all notebooks follow traits discussed in § 2.2, and provide more details on our categorization (and its implications) in the appendix (§ 10).
We also select 146 data science library classes (summarized in LABEL:tbl:library) commonly used in data science notebooks, on which we evaluate Kishu’s correctness and robustness.
@@ Methods
We run notebook cells sequentially and checkpoint after each cell execution. We checkout into the same state for Kishu and Kishu+Det-replay and into a fresh kernel/process for other methods. We measure the (1) checkpoint time (including both tracking and data writing) after each cell execution, (2) checkout time to restore the state from checkpoint files, and (3) tracking overhead spent by Kishu after each cell execution to track updates. We clear the page cache between runs.
@@ 8.2. Notebook Systems and Techniques
There are a variety of works for enhancing data exploration efficiency in notebook-based systems (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Bäuerle et al., 2022; Guo and Seltzer, 2012; Brachmann et al., 2019). Reactive execution engines (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022) track cell reruns and rerun their dependent cells reactively to enforce consistent cell outputs. Notebook recommender systems (Lee et al., 2021; Li et al., 2023b) compute next-cell recommendations based on the current workflow. Symphony (Bäuerle et al., 2022) and B2 (Wu et al., 2020) enable point-and-click interactions with ML models and dataframes, respectively, by translation to equivalent code operations. Burrito (Guo and Seltzer, 2012) and Vizier (Brachmann et al., 2019) construct and visualize a history graph for versioning multi-language (Python, R, SQL) data science code. Diff-in-the-loop (Wang et al., 2022) enables graphical comparisons of dataframes. Kishu facilitates data exploration through efficient time-traveling, and thus, is orthogonal and complementary to these works.

### v4

@@ 5.1. Incremental Checkpointing
The Checkpoint Graph is a directed tree of (incremental) checkpoints representing the branch-based state evolution. Nodes are added for each of Kishu’s checkpoints, and are timestamped with the completion time $t$ of the corresponding cell execution (we refer to the timestamped node and cell execution as node $t$ and CE $t$, respectively). The Checkpoint Graph maintains a head node tracking the current state. Each node $t$ contains the state delta consisting of Co-variables updated by CE $t$. Co-variables stored in each node $t$ are versioned accordingly:
@@
Versioned Co-variables are analogous to versioned datasets: the same Co-variable (w.r.t. variable membership) can take on multiple values during a session being updated by different cell executions. LABEL:fig:checkpoint_graph show an example: CE $t_{3}$ creates the Co-variable {plot}, which is stored in node $t_{3}$ (red) as the Versioned Co-variable $((plot),t_{3})$.
@@
After each CE $t$, Kishu writes a new node $t$ in the Checkpoint Graph with (1) the versioned Co-variables in CE $t$’s state delta, (2) CE $t$’s code, and (3) CE $t$’s accessed versioned Co-variables stored in previous checkpoints (§ 4.3). For example, node $t_{3}$ in LABEL:fig:checkpoint_graph (blue) contains the code (“plot=gmm.result()“) and its dependency on $((gmm),t_{2})$ from node $t_{2}$ (dashed line). Notably, the state delta, code, and variable accesses are respectively analogous to the update, operation, and dependencies in database logging and versioning. The new node $t$ is written under the head node $s$, and and a parent-child relation is added from $s$ to $t$ (now the new head node).
@@
If Kishu cannot write an updated Co-variable into the Checkpoint Graph (e.g., it contains an unserializable object such as a generator (Foundation, 2023b) or hash (Foundation, 2023c)), Kishu simply skips its storage. Instead, upon checkout, the missing Co-variable will be restored through fallback recomputation enabled by the cell code and dependencies stored in the Checkpoint Graph node (§ 5.3).
@@ 5.2. Efficient State Restoration
Kishu’s goal for incremental checkout is to accurately and efficiently restore the current state to the target state. To do so, it must identify the contents of the target state via its timestamp, analogous MVCC’s timestamped snapshots (Bernstein and Goodman, 1983); instead of versioned tables, we identify Versioned Co-variables in the (timestamped) target state:
@@
The Session State at timestamp $t$ is a set of $n$ Versioned Co-variables $\{(\mathcal{X}_{i},t_{i})|1\leq i\leq n\}$ such that for each $(\mathcal{X}_{i},t_{i})$:
@@
    $t_{i}$ is an ancestor of $t$ on the Checkpoint Graph.
@@
    There must not exist another versioned Co-variable $(\mathcal{Y}_{j},t_{j})$ such that $\mathcal{X}_{i}\cup\mathcal{Y}_{j}\neq\emptyset$ and $t_{j}$ is a child of $t_{i}$ and ancestor of $t$.
The session state at timestamp $t$ [^5] Kishu stores snapshots of Session State metadata (i.e., references to contained Co-variables) in Checkpoint Graph nodes. (state $t$ for brevity) is the set of all Versioned Co-variables that are in the namespace after CE $t$, i.e., not overwritten by a newer Versioned Co-variable prior to CE $t$. For example, in LABEL:fig:restore_diff, state $t_{3}$ (top-left) consists of ($(\{plot\},t_{3})$, ($(\{gmm\},t_{2})$, and ($(\{df\},t_{1})$. It does not contain ($(\{gmm\},t_{1})$ as it was overwritten by CE $t_{2}$ (gmm.fit(k=3)) which writes ($(\{gmm\},t_{2})$. Each state $t$ dictates which Versioned Co-variables should be loaded from various Checkpoint Graph nodes for checkouts; for efficient incremental checkout, Kishu identifies the current and target states’ difference w.r.t. the (versioned) Co-variables that need updating: some Co-variables do not need updating when converting the current state to the target state, identifiable via the Checkpoint Graph:
@@
A Co-variable $\mathcal{X}$ is identical between states $t_{a}$ and $t_{b}$ if its versioned counterpart is consistent across $t_{a}$, $t_{b}$, and $t_{c}$, i.e., no CE between (1) nodes $t_{a}$ and $t_{c}$ and (2) nodes $t_{b}$ and $t_{c}$ updated $\mathcal{X}$, hence does not need updating when checking out from $t_{a}$ to $t_{b}$. For example, in LABEL:fig:restore_diff, if checking out from $t_{5}$ to $t_{3}$, the Co-variable {df} (blue) is identical between the states as no CE between (1) nodes $t_{1}$ and $t_{3}$ (2) nodes $t_{1}$ and $t_{5}$ updated it. Otherwise, if the Co-variable $\mathcal{X}$ has diverged between the current and target states, it will need updating (by either loading an appropriate Versioned Co-variable or deleting it) to checkout to the target state. For example, the Co-variable {gmm} (red) has diverged between nodes $t_{5}$ and $t_{3}$ as their parents ($t_{4}$ and $t_{2}$) both updated gmm with their CE (fitting with k=3 and k=10), hence, gmm (and plot) needs updating via loading ({gmm},$t_{2}$) if checking out from state $t_{5}$ to state $t_{3}$.
@@ 5.3. Robust Restoration
As each Checkpoint Graph node $t$ contains the code of CE $t$ and (2) which previous Versioned Co-variables $(\mathcal{X}_{j},t_{j})$ CE $t$ accessed (§ 5.1), any Versioned Co-variable in node $t$’s state delta can be recomputed by (1) loading accessed Versioned Co-variables from previous checkpoints, then re-running CE $t$. For example, in LABEL:fig:restore_fallback, suppose ({plot}, $t_{3}$) (green) fails to load when checking out to state $t_{3}$. (({gmm}, $t_{2}$) is required to rerun CE $t_{3}$ (red); therefore, ({gmm}, $t_{2}$) is loaded from the parent node $t_{2}$, and after rerunning CE $t_{3}$ on the input ({gmm}, $t_{2}$), ({plot}, $t_{3}$) is restored.
@@
Kishu’s fallback recomputation is dynamic and recursive—if another Co-variable is missing or fails to load when retrieving recomputation inputs, fallback recomputation can be recursively performed for that Co-variable. For example, if ({gmm}, $t_{2}$) from node $t_{2}$ fails to load (as fallback recomputation for ({plot}, $t_{3}$) from node $t_{3}$), it itself can be recomputed by loading ({gmm}, $t_{1}$) from node $t_{1}$ and rerunning CE $t_{2}$ (blue).
@@
Kishu guarantees exact restoration for all serializable Co-variables w.r.t. bytestring representation if there are no hidden serialization errors (§ 6.2). While Kishu’s fallback recomputation restores some problematic Co-variables, exactly restoring Co-variables that both (1) fail to store/load and (2) are created non-deterministically (e.g., random generators) is currently unsupported. This limitation is similarly seen in Spark (Zaharia et al., 2010) and Ray’s (Moritz et al., 2018) lineage-based fault tolerance; however, in our case, such objects are rare in data science libraries (§ 7.2), hence we consider this limitation to be acceptable.
@@ 6.1. Implementation
Kishu uses SQLite (SQLite, 2024) to store Versioned Co-variables in the Checkpoint Graph. However, any storage mechanism (Chockchowwat et al., 2022b, a, 2023b) can be used in its place—even in-memory ones if the user wants to maximize checkpointing/checkout efficiency.
@@ 6.2. Design Considerations
Certain object classes may contain incorrect serialization instructions, which, despite being able to be stored/loaded to/from storage, result in silent errors. Kishu currently assumes that instructions are correctly implemented for all objects w.r.t. equality before and after pickling, and does not prevent these silent errors. However, these cases are rare (§ 7.2.1), and Kishu provides a blocklist file for users to force fallback recomputation for Co-variables containing objects belonging to these classes.
@@
While Kishu’s VarGraphs can generalizably detect Co-Variable updates, there are specific cases that allow for more efficient detection methods such as (1) hashing (e.g., XXH64 (xxHash, 2023)) for array-likes (Team, 2024a)) and (2) rule-based static cell (e.g., df.head) identification for skipping update detection (§ 2.2). Kishu currently uses hashing for common array-likes (e.g., tensors (TensorFlow, 2024a)), but can be extended to incorporate (1) other detection methods for specific classes and (2) rule-based cell analyses.
@@ 7. Experimental Evaluation
    Low Checkpoint Times: Kishu’s checkpoints are created up to 5.12$\times$ faster compared to the next best mechanism. (§ 7.4)
@@ 7.1. Experiment Setup
We select 8 data science notebooks from Kaggle Grandmaster-level users or Github-hosted tool tutorials (e.g., Ray) (LABEL:tbl:workload), each featuring a popular data science library, which we categorize as in-progress (3/8) or final (5/8), with the former containing out-of-order cell executions and the latter lacking them. Notably, the final notebooks are also cleaned and contain memory/runtime optimizations. We empirically verify all notebooks follow traits discussed in § 2.2, and provide more categorization details in the appendix (§ 10).
We also select 146 common data science library classes ( LABEL:tbl:library), on which we evaluate Kishu’s correctness and robustness.
@@ Methods
We run notebook cells sequentially from top to bottom and checkpoint after each cell execution. We checkout into the same state for Kishu and Kishu+Det-replay and into a fresh kernel/process for other methods. We measure the (1) checkpoint time (including both tracking and data writing) after each cell execution, (2) checkout time to restore the state from checkpoint files, and (3) tracking overhead of Kishu after each cell execution to track updates. We clear the page cache between runs.
@@ 8.2. Notebook Systems and Techniques
There are a variety of works for enhancing data exploration efficiency in notebook-based systems (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022; Lee et al., 2021; Li et al., 2023b; Bäuerle et al., 2022; Guo and Seltzer, 2012; Brachmann et al., 2019; Chockchowwat et al., 2023a; Hanxi Fang and Park, 2025). Reactive execution engines (Wu et al., 2020; Koop and Patel, 2017; Shankar et al., 2022) track cell reruns and rerun their dependent cells reactively to enforce consistent cell outputs. Notebook recommender systems (Lee et al., 2021; Li et al., 2023b) compute next-cell recommendations based on the current workflow. Symphony (Bäuerle et al., 2022) and B2 (Wu et al., 2020) enable point-and-click interactions with ML models and dataframes, respectively, by translation to equivalent code operations. Burrito (Guo and Seltzer, 2012) and Vizier (Brachmann et al., 2019) construct and visualize a history graph for versioning multi-language (Python, R, SQL) data science code. Diff-in-the-loop (Wang et al., 2022) enables graphical comparisons of dataframes. Kishu facilitates data exploration through efficient time-traveling, and thus, is orthogonal and complementary to these works.